# Scaling laws for post-training quantized large language models

### v1

@@ 2.2 Characteristics of local loss landscape
As the absolute magnitude of ${\bm{w}}$ scales with dimensionality (see Appendix B), we use signal-to-noise ratio (SNR), a relative quantity to measure the magnitude of its perturbation $\Delta{\bm{w}}$,
@@ 2.3 Low-precision data type for quantization
We first ask how the magnitude of quantization errors $\Delta{\bm{w}}=Q({\bm{w}})-{\bm{w}}$ vary across LLMs for certain data types. Despite the existence of significant scaling of $\left\lVert{\bm{w}}\right\rVert$ (see Appendix B for further details), the SQNRs are relatively invariant across model families and model sizes, and vary across numerical data types in a highly predictable manner (see details in Appendix  A). In contrast, NLL losses show a much more nonlinear and less predictable pattern, with a rough trend of lower precision data formats leading to higher losses (see details in Appendix  A).
@@ 6 Limitations
Appendix A SQNR and NLL of MX formats
@@
Appendix B Scaling of $\ell_{2}$-norms of model weights
Figure 9: Scaling of weight $\ell_{2}$-norm. Convention same as in Figure LABEL:fig:loss_scaling. Light gray lines in the background mark square-root power laws, $\left\lVert{\bm{w}}\right\rVert\propto D^{\frac{1}{2}}$.

### v2

@@ 2.2 Characteristics of local loss landscape
As the absolute magnitude of ${\bm{w}}$ scales with dimensionality (see Appendix A), we use signal-to-noise ratio (SNR), a relative quantity to measure the magnitude of its perturbation $\Delta{\bm{w}}$,
@@ 2.3 Low-precision data type for quantization
We first ask how the magnitude of quantization errors $\Delta{\bm{w}}=Q({\bm{w}})-{\bm{w}}$ vary across LLMs for certain data types. Despite the existence of significant scaling of $\left\lVert{\bm{w}}\right\rVert$ (see Appendix A for further details), the SQNRs are relatively invariant across model families and model sizes, and vary across numerical data types in a highly predictable manner (see details in Appendix  B). In contrast, NLL losses show a much more nonlinear and less predictable pattern, with a rough trend of lower precision data formats leading to higher losses (see details in Appendix  B).
@@ 6 Limitations
Appendix A Scaling of $\ell_{2}$-norms of model weights
Figure 7: Scaling of weight $\ell_{2}$-norm. Convention same as in Figure 1 Left. Light gray lines in the background mark square-root power laws, $\left\lVert{\bm{w}}\right\rVert\propto D^{\frac{1}{2}}$.
Appendix B SQNR and NLL of MX formats